If a force of 100N stretches a spring by 0.1cm, find
a) the elastic constant
b) the work done in stretching the spring 0.3cm if the elastic limit is not exceeded
1
Expert's answer
2019-07-03T08:48:35-0400

(a) The Hooke's law states

"F=kx"

So, the elastic constant

"k=\frac{F}{x}=\frac{100\:\rm{N}}{0.001\:\rm{m}}=100000\:\rm{N/m}"

(b) The work done

"W=\frac{1}{2}kx^2"

"=\frac{1}{2}\times 100000\;\rm{N/m}\times (0.003\;\rm{m})^2=0.45\:\rm{J}"